What is a No Cor elation graph?
There are three types of correlation: positive, negative, and none (no correlation).
What is the probability that a sum of 8 on the two dice will occur?
The probability of rolling (2, 6) is (1/6)2=1/36
Add 1/36 for all the other ways to roll 8 - (3,5),(4,4),(5,3),(6,2)
Total probability is 5/36
What is the probability of throwing a 3 or 6 if you role two dice at once?
In this case, you should first calculate the probability of NOT rolling either a 3 or a 6. For one die, this probability is 4/6 = 2/3. For two dice, this probability is 2/3 x 2/3 = 4/9. Therefore, the probability that you DO get a 3 or a 6 is 1 - 4/9 = 5/9.
What are the structures of a stem?
On a stem, there are nodes, which is the place where the leaf attaches to the stem. There are also leaf scars, which are places where old leaves used to be attached. There are the internodes, which are the areas between the nodes. The lateral buds are located at the nodes, and they are underdeveloped tissues that turn into the branches and leaves that we see now. Finally, there are terminal buds, which are the reason trees are able to grow upwards. Once the bud develops, it leaves a bud scale scar, which is symbolized by rings around the tree branch.
What it the sample space for the number of heads than can appear if a coin is rolled three times?
Each roll of the coin can either have a head or not have a head. So in three rolls there could be:
giving a sample space for the number of heads in three rolls of a coin:
{0, 1, 2, 3}
What is the least even number having six digits?
The smallest six-digit number is 100000 (one hundred thousand), which is also an even number.
Pl tell the a to z numaric value?
what? i am just gonna tell you pi=3.14 cause i dont get what you mean.
1 From a pack of 52 playing cards what is the probability of picking?
there are 4 1's or aces in a pack of 52 cards. Probability of picking an ace=4/52=1/13
1) When a company is charged with discrimination (e.g. racial or gender), the plaintiff must justify the claim. They typically quote the number of jobs filled and how many of each group (men vs women, white vs black, etc.) were awarded those jobs. Their aim is to show that a certain group (e.g. men or whites) got more of the jobs than they should have. Assuming women were 50% of applicants for a job but men were hired 75% of the time, they might use that as a (statistical) basis for a claim of discrimination. In all fairness, there could be many other reasons for a hiring or promotional bias, but that is beyond the scope of this answer.
2) Bias in application of the death penalty. There have been many claims that black males are given the death penalty more easily than others. Again, this is easily disputable, but that is beyond this scope.
3) Pollution. If cancer rates are (statistically) high around a chemical plant or fracking operation, this can be the basis of a law suit.
4) Education. Claims of "failing schools" are based on comparing student test scores.
5) DNA as evidence. To claim a match in DNA found at a crime scene vs the accused, the prosecutor explains that the chance of anyone matching is, for example, 1 in 100 billion, therefore the accused must be the one who committed the crime.
6) Options fixing. Many company executives were caught getting stock options on days when the company's stock price was surprisingly low based on its running average. By statistics, the prosecutor was able to show that the chance of this happening was vanishingly low, therefore the assignment date must have been chosen retroactively.
7) Accounting Fraud. Certain types of incomes have certain distributions, e.g. dollar amounts start with 1 (e.g. $1.50, $10.23) much more often than they start with 9 (e.g. $9.50, $9.23). So a non-standard distribution of dollar amounts in a database is a strong indicator of fraud.
The answer can be arrived at by using a compound annual growth rate (CAGR) calculator. In this case an initial deposit of $2,000 growing to $4,706 after seven years would equate to a 13 percent annual rate of return.
What is the least common multiple of 14 and 31?
the least common multiple of these two values is 434 (their product).
Find the probability of rolling a 10 with two dice?
Assuming that the random variable is the sum of the two numbers rolled, the answer is 3/36 or 1/12.
It means that you are a very average student.
Are m and f mutually exclusive?
If by this you mean male and female then no. Although rare individuals with ambiguous characteristics do occur naturally in human and animal populations. Please see the wikipedia article entitled 'intersex' for more information.
What is the value of z in z 1755?
Unfortunately, limitations of the browser used by Answers.com means that we cannot see most symbols. It is therefore impossible to give a proper answer to your question. Please resubmit your question spelling out the symbols as "plus", "minus", "times", "equals". There is no equation (nor inequality) in the question.
What is the standard normal deviation?
The standard deviation of a normal deviation is the square root of the mean, also the square root of the variance.